# Improving Assembly Code Performance with Large Language Models via Reinforcement Learning

### v1

@@ Abstract
Large language models (LLMs) have demonstrated strong performance across a wide range of programming tasks, yet their potential for code optimization remains underexplored. This work investigates whether LLMs can optimize the performance of assembly code, where fine-grained control over execution enables improvements that are difficult to express in high-level languages. We present a reinforcement learning framework that trains LLMs using Proximal Policy Optimization (PPO), guided by a reward function that considers both functional correctness, validated through test cases, and execution performance relative to the industry-standard compiler gcc -O3. To support this study, we introduce a benchmark of 8,072 real-world programs. Our model, Qwen2.5-Coder-7B-PPO, achieves 96.0% test pass rates and an average speedup of 1.47× over the gcc -O3 baseline, outperforming all 20 other models evaluated, including Claude-3.7-sonnet. These results indicate that reinforcement learning can unlock the potential of LLMs to serve as effective optimizers for assembly code performance.
@@ 1 Introduction
Figure 1: Overview of the assembly code optimization task. Given a C program and its baseline assembly from gcc -O3, an LLM is fine-tuned with Proximal Policy Optimization (PPO) to generate improved assembly. The reward function reflects correctness and performance based on test execution.
Recent advances in large language models (LLMs) have achieved state-of-the-art solutions across a wide range of programming tasks [1, 2, 3, 4, 5]. However, their potential for program optimization remains underexplored. Generating highly optimized code is critical in performance-sensitive domains, and prior work has investigated the use of LLMs to optimize C++ and Python programs [6, 7, 8]. In this work, we aim to utilize LLMs to improve the performance of assembly code, extending their capabilities beyond optimization for high-level languages.
Assembly code optimization is traditionally the responsibility of compilers. While modern compilers apply a series of rule-based transformations to improve performance, such a design introduces the classic phase ordering problem [9], where the order of optimizations can substantially affect the performance of the generated code. Due to the inherent complexity of the optimization task, especially the vast space of possible transformation sequences, compilers face fundamental challenges in converging to optimal code, often leaving significant performance on the table [10].
An alternative approach is superoptimization, which searches the space of all programs that are functionally equivalent to the compiler’s output, aiming to identify the most performant variant. In principle, this strategy may yield optimal code. However, the search space grows exponentially with program size, making exhaustive exploration computationally infeasible in practice. Furthermore, prior work on superoptimization [11, 12] has primarily targeted loop-free, straight-line code, where it is more tractable to formally verify that the optimized code is semantically equivalent to the original. As a result, these approaches are not directly applicable to most real-world programs with loops.
In this work, we explore using large language models to optimize the performance of assembly code. Compared with high-level languages such as Python or C++, assembly code operates closer to the hardware, offering fine-grained control over execution and enabling optimizations that are difficult to express or realize in higher-level code. However, this setting poses several challenges. Assembly code is relatively rare and may be underrepresented in pretraining corpora [13], making it harder for LLMs to reason effectively about their behavior. Furthermore, industry compilers such as GCC have been extensively tuned by performance engineers over decades. Achieving additional speedups beyond gcc -O3 (the compiler’s highest optimization level) is a technically challenging task.
To address the challenges of low-level code optimization, we apply reinforcement learning to enhance the ability of LLMs to optimize assembly code. As shown in Figure 1, we use Proximal Policy Optimization (PPO) to train an LLM using a reward function that considers both correctness and performance. Correctness is evaluated based on whether the generated code passes program-specific test cases, and performance is measured by its speedup relative to the baseline produced by gcc -O3. To support this setting, we construct a new dataset of 8,072 assembly programs derived from real-world competitive programming submissions. Each instance includes input-output test cases and baseline assembly code generated by the compiler at the highest optimization level, which serves as the starting point for further optimization.
We evaluate our approach on the proposed benchmark and find that reinforcement learning substantially improves the ability of LLMs to optimize assembly code. Starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves a modest 1.10x speedup over the gcc -O3 baseline, our PPO-trained model reaches 1.47× average speedup and improves both compile and test pass rates to 96.0%. It achieves the strongest performance across all evaluation metrics, outperforming all 20 other models evaluated, including Claude-3.7-sonnet. Ablation studies show that reward functions emphasizing final speedup, rather than intermediate correctness signals, lead to more effective training.
@@
  • We introduce the task of optimizing assembly code performance using large language models, aiming for fine-grained performance improvements beyond what traditional compiler optimizations can achieve.
  • We construct a dataset of real-world C programs paired with the corresponding assembly code generated by the gcc -O3 baseline. Using this dataset, we explore improving LLMs on this task through reinforcement learning, applying Proximal Policy Optimization (PPO).
  • We evaluate 21 LLMs on the proposed benchmark and show that our training substantially improves performance: Qwen2.5-Coder-7B-PPO achieves the highest compile and test pass rates, as well as the best average speedup (1.47×) over the gcc -O3 baseline, outperforming all other models (including Claude-3.7-sonnet) across all evaluation metrics.
@@ 2 Related Work
Benchmarks for evaluating large language models (LLMs) on code generation from natural language specifications have received increasing attention. Notable examples include HumanEval [1], MBPP [2], APPS [3], and more recent efforts [14, 15, 16, 17]. In parallel, many models have been developed to enhance code generation capabilities, such as Codex [1], AlphaCode [18], CodeGen [19], InCoder [20], StarCoder [21], DeepSeek-Coder [22], Code Llama [23], and others [24, 25]. Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair [26, 27], software testing [28, 29], bug localization [30], and transpilation [31, 32]. SWE-bench [4] integrates these tasks into a benchmark for resolving real GitHub issues. Building on this, SWE-agent [5] and subsequent works [33, 34] employ an agent-based framework that leverages LLMs to improve the issue resolution process.
Recent work has also explored LLMs for improving program performance. CodeRosetta [35] targets automatic parallelization, such as translating C++ to CUDA. Other efforts focus on optimizing Python code for efficiency [7, 8] or enabling self-adaptation [36], and improving C++ performance [6]. Of particular relevance are approaches to low-level code optimization [37, 38]. The LLM Compiler foundation models [39, 40] are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance. LLM-Vectorizer [41] offers a formally verified solution for auto-vectorization, a specific compiler pass. In contrast, our work does not restrict the optimization type and uses test-case validation.
@@
The space of code optimization is vast, and many approaches have leveraged machine learning to improve program performance. A classic challenge in compilers is the phase-ordering problem, where performance depends heavily on the sequence of optimization passes. AutoPhase [42] uses deep reinforcement learning to tackle this, while Coreset [43] employs graph neural networks (GNNs) to guide optimization decisions. Modern compilers apply extensive rewrite rules but offer no guarantee of optimality. Superoptimization seeks the most efficient program among all semantically equivalent variants of the compiler output. Traditional methods use stochastic search, such as Markov Chain Monte Carlo [11], with follow-up work improving scalability [44, 12] and extending to broader domains [45, 46]. These rely on formal verification for correctness, restricting them to small, loop-free programs. In contrast, our approach uses test-based validation, enabling optimization of general programs with loops. With the rise of deep learning, substantial attention has turned to optimizing GPU kernel code. AutoTVM [47] pioneered statistical cost model-based search for CUDA code optimization, followed by methods such as Ansor [48], AMOS [49], and other recent systems [50, 51, 52].
More recently, using LLMs as code optimizers has gained popularity [6, 53, 37], with growing attention to reinforcement learning approaches that guide LLMs through reward-based feedback [54, 34]. CodeRL [55] incorporates unit test-based rewards within an actor-critic framework [56], while PPOCoder [57] extends this with Proximal Policy Optimization (PPO) [58], along with other variants [59]. Subsequent efforts have adapted RL-based techniques [60] to additional low-resource programming languages, including Verilog [61]. To the best of our knowledge, our work is the first to apply reinforcement learning to optimize assembly code using LLMs. Assembly code offers fine-grained control and potential for significant performance gains, but it remains underexplored due to limited training data and the complexity of low-level semantics.
@@ 3.1 Task Definition
Let $C$ be a program written in a high-level language such as C. A modern compiler like gcc can compile $C$ into an x86-64 assembly program $P=\texttt{gcc}(C)$, which can then be further assembled into an executable binary. The assembly program $P$ serves as an intermediate representation that exposes low-level optimization opportunities, making it suitable for aggressive performance improvement. We assume the semantics-preserving nature of the compilation process, i.e., $\llbracket C\rrbracket=\llbracket P\rrbracket$, so that the behavior of the assembly program $P$ is identical to that of the source program $C$.
@@
We say that an assembly program $P^{\prime}$ is valid if it can be successfully assembled and linked into an executable binary. Let $\texttt{valid}(P^{\prime})\in\{\texttt{True},\texttt{False}\}$ denote this property. We define the set of correct programs as:
@@
Let $t(P)$ denote the execution time of $P$ on the test set $\mathcal{T}$, and let $t(P^{\prime})$ be the corresponding execution time for $P^{\prime}$. The speedup of $P^{\prime}$ relative to $P$ is defined as:
@@
The objective is to generate a candidate program $P^{\prime}$ that maximizes $\text{Speedup}(P^{\prime})$. Only programs in $\mathcal{S}(P)$ are eligible for speedup; any candidate that fails to compile into a binary or produces incorrect outputs is assigned a default speedup of 1. This reflects a practical fallback: when the generated program is invalid, the system can revert to the baseline $P$, compiled with gcc -O3, which defines the 1× reference performance. Although $\mathcal{S}(P)$ captures the correctness criteria, we do not restrict the LLM to generate only valid programs. Instead, the model produces arbitrary assembly code, and correctness is verified post hoc via compilation and test execution. We train an LLM using reinforcement learning (see Section 3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
@@ 3.2 Dataset Construction
We construct our dataset using C programs from CodeNet [62], a large-scale corpus of competitive programming submissions. Each dataset instance is a tuple $(C,P,\mathcal{T})$, where $C$ is the original C source code, $P=\texttt{gcc\_O3}(C)$ is the corresponding x86-64 assembly generated by compiling $C$ with gcc at the -O3 optimization level, and $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is the test set. Since not all CodeNet problems include test inputs, we adopt those provided by prior work [18] to define $x_{i}$, but discard their output labels. Instead, we regenerate each output $y_{i}$ by executing the original submission on input $x_{i}$, as many CodeNet programs are not accepted solutions, and even accepted ones do not reliably pass all test cases.
Given the scale of CodeNet, which contains over 8 million C and C++ submissions, we sample a subset for this study. To focus on performance-critical cases, we sample programs that exhibit the highest relative speedup from gcc -O0 (no optimization) to gcc -O3 (maximum optimization). Such strategy serves two purposes: (1) it favors programs with complex logic that lead to suboptimal performance under -O0 and can be effectively optimized by -O3, and (2) it creates a more challenging setting by starting from code that has already benefited from aggressive compiler optimizations. If an LLM can generate code that further improves upon gcc -O3, it suggests that the model can outperform the compiler’s “expert” solution. The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section 4.
@@ 3.3 Reinforcement Learning
We conceptualize our task as a standard contextual multi-armed bandit problem [63], defined by a context space $\mathcal{S}$, an action space $\mathcal{A}$, and a reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Each context $s\in\mathcal{S}$ represents a problem instance, comprising the source program $C$, its baseline assembly $P$, and the associated test cases $\mathcal{T}$. An action $a\in\mathcal{A}$ corresponds to generating a candidate assembly program $\tilde{P}$. The reward function $r(s,a)$ evaluates the quality of the generated program based on correctness and performance. We will describe different designs of the reward function later. A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ maps a context $s$ to a probability distribution over actions and samples an action $a\in\mathcal{A}$ stochastically. Given a distribution $\mu$ over problem instances, the expected performance of a policy $\pi$ under reward function $r$ is expressed as $\mathbb{E}_{s\sim\mu,a\sim\pi(\cdot\mid s)}\left[r(s,a)\right]$. The objective is to find a policy that maximizes this expected reward.
Optimization with PPO.
We train the policy using Proximal Policy Optimization (PPO) [58], a first-order policy-gradient algorithm that stabilizes training by constraining each policy update to remain close to the previous one. Specifically, PPO maximizes a clipped surrogate objective of the form $\mathbb{E}_{s,a}\left[\min\left(\rho(\theta)\hat{A},\;\operatorname{clip}(\rho(\theta),1-\epsilon,1+\epsilon)\,\hat{A}\right)\right],$ where $\rho(\theta)=\pi_{\theta}(a\mid s)/\pi_{\theta_{\text{old}}}(a\mid s)$ is the probability ratio between the current and previous policy, $\hat{A}$ is the estimated advantage of action $a$ in state $s$, and $\epsilon$ is a clipping coefficient that limits the policy update to a small trust region. We use a critic model to estimate $\hat{A}$, and compute rewards based on the correctness and execution time of the generated program, eliminating the need for a separate reward model.
@@
$\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],\quad\operatorname{speedup}(s,a)=t(P)/t(\tilde{P}),$
which respectively denote the fraction of test cases passed and the speedup of the generated program $\tilde{P}$ relative to the baseline $P$. We evaluate two reward function variants:
  1. 1.
    Correctness-Guided Speedup (CGS):
    $r(s,a)=\begin{cases}-1,&\text{if }\tilde{P}\text{ fails to compile},\\[4.0pt] \operatorname{pass}(s,a),&\text{if some tests fail},\\[4.0pt] 1+\alpha\cdot\operatorname{speedup}(s,a),&\text{if all tests pass}.\end{cases}$
  2. 2.
    Speedup-Only (SO):
    $r(s,a)=\begin{cases}0,&\text{if }\tilde{P}\text{ fails to compile or any test fails},\\[4.0pt] \operatorname{speedup}(s,a),&\text{otherwise}.\end{cases}$
In CGS, the constant $\alpha$ controls the relative importance of speedup once full correctness is achieved (i.e., all test cases pass). The CGS reward provides a dense signal by assigning intermediate credit for successful compilation and partially correct outputs, guiding the policy even when the final objective is not yet met. In contrast, SO defines a more direct and sparse objective, assigning nonzero reward only to programs that are both correct and performant, thereby rewarding only the terminal goal of achieving speedup.

### v2

# SuperCoder: Assembly Program Superoptimization with Large Language Models

@@ Abstract
Superoptimization is the task of transforming a program into a faster one while preserving its input–output behavior. In this work, we investigate whether large language models (LLMs) can serve as superoptimizers, generating assembly programs that outperform code already optimized by industry-standard compilers. We construct the first large-scale benchmark for this problem, consisting of 8,072 assembly programs averaging 130 lines, in contrast to prior datasets restricted to 2–15 straight-line, loop-free programs. We evaluate 23 LLMs on this benchmark and find that the strongest baseline, Claude-opus-4, achieves a 51.5% test-passing rate and a 1.43× average speedup over gcc -O3. To further enhance performance, we fine-tune models with reinforcement learning, optimizing a reward function that integrates correctness and performance speedup. Starting from Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup), the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup. Our results demonstrate, for the first time, that LLMs can be applied as superoptimizers for assembly programs, establishing a foundation for future research in program performance optimization beyond compiler heuristics.
@@ 1 Introduction
Superoptimization is the task of transforming a program into a faster one while preserving its input-output behavior. In this work, we investigate whether large language models (LLMs) can perform superoptimization by generating assembly code that surpasses the performance of compiler outputs.
Decades of research have tackled the problem of code optimization, giving rise to two main approaches. The first develops better algorithms for rule-based transformations in compilers (Wolf & Lam, 1991). However, given the vast space of possible transformations, compiler-optimized code is not guaranteed to be optimal and often leaves performance untapped (Center et al., 1971; Theodoridis et al., 2022). The second, superoptimization, develops search algorithms that directly explore the space of all possible programs, aiming to discover the correct variant with the best performance rather than relying on a fixed set of transformation rules (Schkufza et al., 2013).
Superoptimization is an aggressive form of program optimization that can outperform compiler-optimized code, yet existing literature has focused on very short, straight-line assembly programs without loops. Prior work has primarily relied on CPU-based search heuristics, which fail to scale to larger programs (Schkufza et al., 2013; Phothilimthana et al., 2016; Koenig et al., 2021b); available datasets include at most 15 lines of straight-line assembly (Koenig et al., 2021a).
@@
Figure 1: Overview of the assembly code optimization task. Given a C program and its baseline assembly from gcc -O3, an LLM is fine-tuned with PPO or GRPO to generate improved assembly. The reward function reflects correctness and performance based on test execution.
In this work, we explore the use of LLMs as a superoptimizer to improve the performance of assembly code. In contrast to most prior work on code generation from natural language (Chen et al., 2021; Austin et al., 2021; Hendrycks et al., 2021; Zhuo et al., 2024), we tackle a fundamentally different and more technically demanding task: improving assembly code that has already been optimized by the industry-standard compiler at its highest optimization level (gcc -O3). Compilers have been refined over decades of expert-driven development, and surpassing them remains a central challenge in programming languages, as compilers form the foundation of all software.
Unlike high-level programming languages (e.g., Python or C), large-scale, high-quality assembly datasets are scarce. As the first study in this direction, we construct a dataset of 8,072 assembly programs. Each instance includes input–output test cases and baseline assembly generated by the compiler at its highest optimization level (gcc -O3), which serves as the starting point for further optimization. In contrast, the datasets commonly used in the superoptimization community (Warren, 2013; Schkufza et al., 2013; Koenig et al., 2021a) are either extremely limited in size, containing only 25 programs, or consist of toy examples with 2 to 15 instructions, 74% of which have between 3 and 6 instructions and contain no loops. Our dataset is substantially larger, with assembly programs averaging 130 lines and including loops. Our test suites achieve 96.2% line and 87.3% branch coverage, demonstrating strong test quality. Our dataset represents a substantial step forward in scale for evaluating superoptimization techniques.
Beyond evaluating existing models, we also apply reinforcement learning (RL) for fine-tuning to further enhance their capabilities. We use widely adopted algorithms, including Proximal Policy Optimization (PPO) and Group Relative Policy Optimization (GRPO), to train an LLM with a reward function that integrates both correctness and performance speedup. Prior work on LLM-based performance optimization has explored alternative methodologies such as supervised fine-tuning (Shypula et al., 2023), chain-of-thought prompting (Liu et al., 2024c), agent-based frameworks (Wei et al., 2025b; c), and preference learning (Du et al., 2024). Our approach optimizes speedup explicitly in the reward function, making reinforcement learning well suited to superoptimization. To our knowledge, this is the first application of reward-based RL to LLMs for code performance optimization, with correctness and speedup jointly encoded in the objective.
We evaluate 23 LLMs on this task and find that the best-performing model, Claude-opus-4, achieves a 51.5% test-passing rate and an average speedup of 1.43× over the compiler-optimized baseline (gcc -O3). Our reinforcement learning approach is highly effective: starting from the base model Qwen2.5-Coder-7B-Instruct, which achieves 61.4% correctness and a modest 1.10× speedup, the fine-tuned model SuperCoder attains 95.0% correctness and 1.46× average speedup. We further analyze the code transformation patterns learned by the model.
@@
  • We are the first to introduce superoptimization as a task for LLMs, a technically demanding challenge that aims to improve assembly code already optimized by industry-standard compilers.
  • We construct the first large-scale dataset of 8,072 assembly programs, averaging 130 lines. This far surpasses prior loop-free datasets under 15 lines and marks a substantial step forward in scale and realism for evaluating superoptimization techniques.
  • We evaluate 23 LLMs on the benchmark and show that RL-based training substantially improves performance: fine-tuning Qwen2.5-Coder-7B-Instruct (61.4% correctness, 1.10× speedup) results in SuperCoder with 95.0% correctness and 1.46× speedup.
@@ 2 Related Work
Benchmarks for evaluating large language models (LLMs) on code generation from natural language specifications have received increasing attention. Notable examples include HumanEval (Chen et al., 2021), MBPP (Austin et al., 2021), APPS (Hendrycks et al., 2021), and more recent efforts (Liu et al., 2023b; Li et al., 2024; Xia et al., 2024b; Zhuo et al., 2024). In parallel, many models have been developed to enhance code generation capabilities, such as Codex (Chen et al., 2021), AlphaCode (Li et al., 2022), CodeGen (Nijkamp et al., 2022), InCoder (Fried et al., 2022), StarCoder (Li et al., 2023), DeepSeek-Coder (Guo et al., 2024), Code Llama (Roziere et al., 2023), and others (Hui et al., 2024; Wei et al., 2025e). Beyond code generation, LLMs have been applied to real-world software engineering tasks including automated program repair (Xia & Zhang, 2022; Xia et al., 2023b), software testing (Xia et al., 2023a; Deng et al., 2024), bug localization (Yang et al., 2024a), transpilation (Yang et al., 2024c; Bhatia et al., 2024), equivalence checking (Wei et al., 2025a), and synthesis (Wei et al., 2025d). SWE-bench (Jimenez et al., 2023) integrates these tasks into a benchmark for resolving real GitHub issues. Building on this, SWE-agent (Yang et al., 2024b) and subsequent works (Xia et al., 2024a; Wei et al., 2025f) employ an agent-based framework that leverages LLMs to improve the issue resolution process.
Recent work has also explored LLMs for improving program performance. CodeRosetta (TehraniJamsaz et al., 2024) targets automatic parallelization, such as translating C++ to CUDA. Other efforts focus on optimizing Python code for efficiency (Du et al., 2024; Liu et al., 2024c) or enabling self-adaptation (Huang et al., 2024), and improving C++ performance (Shypula et al., 2023). Of particular relevance are approaches to low-level code optimization (Wei et al., 2024; Ouyang et al., 2025). The LLM Compiler foundation models (Cummins et al., 2024; 2025) are primarily designed for code size reduction and binary disassembly, whereas our work focuses on optimizing assembly code for performance. LLM-Vectorizer (Taneja et al., 2025) offers a formally verified solution for auto-vectorization, a specific compiler pass. In contrast, our work does not restrict the optimization type and uses test-case validation.
@@
The space of code optimization is vast, and many approaches have leveraged machine learning to improve program performance. A classic challenge in compilers is the phase-ordering problem, where performance depends heavily on the sequence of optimization passes. AutoPhase (Haj-Ali et al., 2020) uses deep reinforcement learning to tackle this, while Coreset (Liang et al., 2023) employs graph neural networks (GNNs) to guide optimization decisions. Modern compilers apply extensive rewrite rules but offer no guarantee of optimality. Superoptimization seeks the most efficient program among all semantically equivalent variants of the compiler output. Traditional methods use stochastic search, such as MCMC sampling (Schkufza et al., 2013), with follow-up work improving scalability (Phothilimthana et al., 2016; Bunel et al., 2016) and extending to broader domains (Sharma et al., 2015; Churchill et al., 2017). These rely on formal verification for correctness, restricting them to small, loop-free programs. In contrast, our approach uses test-based validation, enabling optimization of general programs with loops. With the rise of deep learning, substantial attention has turned to optimizing GPU kernel code. AutoTVM (Chen et al., 2018) pioneered statistical cost model-based search for CUDA code optimization, followed by methods such as Ansor (Zheng et al., 2020), AMOS (Zheng et al., 2022), and others (Shao et al., 2022; Zhao et al., 2024; Wu et al., 2024).
More recently, LLMs have been explored as code optimizers (Shypula et al., 2023; Grubisic et al., 2024; Wei et al., 2024; 2025b; 2025c), with growing interest in reinforcement learning that guides generation through reward signals (Dou et al., 2024; Wei et al., 2025f). Rewards are often derived from unit-test correctness (Le et al., 2022; Shojaee et al., 2023; Liu et al., 2023a) or binary preference signals (Liu et al., 2024b; Du et al., 2024). To our knowledge, this is the first work to apply reinforcement learning to optimize code performance with LLMs, with concurrent efforts exploring CUDA kernel optimization (Li et al., 2025; Baronio et al., 2025).
@@ 3.1 Task Definition
Let $C$ be a program written in a high-level language such as C. A modern compiler like gcc can compile $C$ into an x86-64 assembly program $P=gcc(C)$, which can then be further assembled into an executable binary. The assembly program $P$ serves as an intermediate representation that exposes low-level optimization opportunities, making it suitable for aggressive performance improvement. We assume the semantics-preserving nature of the compilation process, i.e., $\llbracket C\rrbracket=\llbracket P\rrbracket$, so that the behavior of the assembly program $P$ is identical to that of the source program $C$.
@@
We say that an assembly program $P^{\prime}$ is valid if it can be successfully assembled and linked into an executable binary. Let $\texttt{valid}(P^{\prime})\in\{\texttt{True},\texttt{False}\}$ denote this property. Based on all that we said above, we define the set of correct programs as:
@@
Let $t(P)$ denote the execution time of $P$ on the test set $\mathcal{T}$, and let $t(P^{\prime})$ be the corresponding execution time for $P^{\prime}$. The speedup of $P^{\prime}$ relative to $P$ is defined as follows. If the LLM-generated program is invalid or slower, we fall back to the baseline and assign a speedup of 1.
@@
The objective is to generate a candidate program $P^{\prime}$ that maximizes $\text{Speedup}(P^{\prime})$. Only programs in $\mathcal{S}(P)$ are eligible for speedup; any candidate that fails to compile into a binary or produces incorrect outputs is assigned a default speedup of 1. This reflects a practical fallback: when the generated program is invalid, the system can revert to the baseline $P$, compiled with gcc -O3, which defines the 1× reference performance. Although $\mathcal{S}(P)$ captures the correctness criteria, we do not restrict the LLM to generate only valid programs. Instead, the model produces arbitrary assembly code, and correctness is validated post hoc via compilation and test execution. We train an LLM using reinforcement learning (see Section˜3.3) to generate candidates that both satisfy correctness and achieve performance improvements.
@@ 3.2 Dataset Construction
We construct our dataset using C programs from CodeNet (Puri et al., 2021), a large-scale corpus of competitive programming submissions. Each dataset instance is a tuple $(C,P,\mathcal{T})$, where $C$ is the original C source code, $P=gcc(C)$ is the corresponding x86-64 assembly generated by compiling $C$ with gcc at the -O3 optimization level, and $\mathcal{T}=\{(x_{i},y_{i})\}_{i=1}^{n}$ is the test set. Since not all CodeNet problems include test inputs, we adopt those provided by prior work (Li et al., 2022) to define $x_{i}$, but discard their output labels. Instead, we regenerate each output $y_{i}$ by executing the original submission on input $x_{i}$, as many CodeNet programs are not accepted solutions, and even accepted ones do not reliably pass all test cases.
Given the scale of CodeNet, which contains over 8 million C and C++ submissions, we sample a subset for this study. To focus on performance-critical cases, we sample programs that exhibit the highest relative speedup from gcc -O0 (no optimization) to gcc -O3 (maximum optimization). Such a strategy serves two purposes: (1) it favors programs with complex logic that lead to suboptimal performance under -O0 and can be effectively optimized by -O3, and (2) it creates a more challenging setting by starting from code that has already benefited from aggressive compiler optimizations. If an LLM can generate code that further improves upon gcc -O3, it suggests that the model can outperform the compiler’s “expert” solution. The final dataset consists of 7,872 training programs and 200 held-out evaluation programs, with additional statistics provided in Section˜4.
@@ 3.3 Reinforcement Learning
We conceptualize our task as a standard contextual multi-armed bandit problem (Lu et al., 2010), defined by a context space $\mathcal{S}$, an action space $\mathcal{A}$, and a reward function $r:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$. Each context $s\in\mathcal{S}$ represents a problem instance, comprising the source program $C$, its baseline assembly $P$, and the associated test cases $\mathcal{T}$. An action $a\in\mathcal{A}$ corresponds to generating a candidate assembly program $\tilde{P}$. The reward function $r(s,a)$ evaluates the quality of the generated program based on correctness and performance. We will describe different designs of the reward function later. A policy $\pi:\mathcal{S}\rightarrow\Delta(\mathcal{A})$ maps a context $s$ to a probability distribution over actions and samples an action $a\in\mathcal{A}$ stochastically. Given a distribution $\mu$ over problem instances, the expected performance of a policy $\pi$ under reward function $r$ is expressed as $\mathbb{E}_{s\sim\mu,a\sim\pi(\cdot\mid s)}\left[r(s,a)\right]$. The objective is to find a policy that maximizes this expected reward.
Optimization with PPO and GRPO.
[…]
@@
$t(P~),\operatorname{pass}(s,a)=\tfrac{1}{|\mathcal{T}|}\sum_{(x,y)\in\mathcal{T}}\mathbf{1}[\tilde{P}(x)=y],\quad\operatorname{speedup}(s,a)$$t(\tilde{P})$,
which denote the fraction of test cases passed and the speedup of the generated program $\tilde{P}$ relative to the baseline $P$. We use the following reward function during training:
$r(s,a)=\begin{cases}0,&\text{if }\operatorname{pass}(s,a)<1,\\[4.0pt] \operatorname{speedup}(s,a),&\text{if }\operatorname{pass}(s,a)=1.\end{cases}$
If a generated program fails to compile or does not pass all tests, its reward is set to 0, with no partial credit for partial correctness. Only when the code compiles and passes all tests is a positive reward assigned, equal to the achieved speedup.